# arXiv:2602.23696v1[cs.LG] 27 Feb 2026

### v1

@@ 4.4 Block Localization: The First Transformer Block Drives the Backbone
Checkpoint Blk 0 Blk 1 Blks 2–5 Blk 6 Blk 7
Step 200 (init) 0.014 0.003 0.001 0.000 0.002
Step 1800 (peak) 0.019 0.001 0.001 0.003 0.004
Step 2000 (trough) 0.029 0.001 0.001 0.005 0.005
[…]
The cumulative signed bias, obtained by summing $b_{\ell}(t)$ across checkpoints,
$B_{\ell}\;=\;\sum_{t\in\mathcal{T}}b_{\ell}(t),$ (25)
confirms this: Block 0 accumulates monotonically to $B_{0}\approx 0.12$, while blocks 1–7 remain near 0.01. The same pattern holds in seed 271.
Interpretation. Block 0 (the first transformer block) sits directly above the token embeddings and is the first to process the LM-relevant token representations. Its attention and MLP weight matrices are the most directly constrained by the language modeling objective, giving the LM gradient at this level the highest temporal coherence (Saxe et al., 2014)—it consistently pulls these parameters in a stable direction determined by the language statistics.
4.5 Fisher Curvature: The Backbone Stiffens
Does the backbone become progressively harder to move along? We answer this using the Fisher information matrix, which characterizes the local curvature of the loss landscape.
Definition 3 (Empirical Fisher and Rayleigh quotient).
Given $M$ mini-batch gradients $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{M}\}$ stacked into a matrix $\mathbf{G}\in\mathbb{R}^{M\times D}$, the empirical Fisher is
$\hat{\mathbf{F}}\;=\;\frac{1}{M}\,\mathbf{G}^{\!\top}\mathbf{G}\;=\;\frac{1}{M}\sum_{i=1}^{M}\mathbf{g}_{i}\,\mathbf{g}_{i}^{\!\top}.$ (26)
The Rayleigh quotient of a unit vector $\mathbf{v}$ with respect to $\hat{\mathbf{F}}$ is
$q(\mathbf{v})\;=\;\mathbf{v}^{\!\top}\hat{\mathbf{F}}\,\mathbf{v}\;=\;\frac{1}{M}\,\lVert\mathbf{G}\mathbf{v}\rVert^{2}\;=\;\frac{1}{M}\sum_{i=1}^{M}\langle\mathbf{g}_{i},\,\mathbf{v}\rangle^{2}.$ (27)
Remark 2 (Computational trick).
Computing eq.˜27 requires only a single matrix-vector product $\mathbf{G}\mathbf{v}\in\mathbb{R}^{M}$, avoiding construction of the $D\times D$ Fisher matrix. Since $D\approx 25\times 10^{6}$ (full trunk) and $M=32$, this reduces memory from $O(D^{2})$ to $O(MD)$.
We define the anisotropy ratio as the Rayleigh quotient along the backbone relative to the average over random orthogonal directions:
Definition 4 (Anisotropy ratio).
$\alpha\;=\;\frac{q(\mathbf{v}_{\mathrm{b}})}{\frac{1}{K}\sum_{k=1}^{K}q(\mathbf{w}_{k})},$ (28)
where $\{\mathbf{w}_{1},\ldots,\mathbf{w}_{K}\}$ are $K=10$ random unit vectors orthogonal to $\mathbf{v}_{\mathrm{b}}$, generated by Gram–Schmidt orthogonalization of random Gaussian vectors. $\alpha>1$ indicates that the loss landscape is stiffer along the backbone than along typical directions.
Table 6: Rayleigh quotients and anisotropy (seed 42). $M=32$ mini-batches.
Step $q(\mathbf{v}_{\mathrm{b}})$ $q(\mathbf{v}_{\mathrm{sw}})$ $q(\mathbf{v}_{\mathrm{PC2}})$ Anisotropy $\alpha$
200 (init) $2.4\times 10^{-6}$ $2.9\times 10^{-6}$ $3.2\times 10^{-6}$ $1.3\times$
1800 (peak) $5\times 10^{-6} 2.2$$2\times 10^{-6}$ $5.2\times 10^{-6}$ $1.9\times$
4800 (transition) $\mathbf{1.6\times 10^{-4}}$ $8.2\times 10^{-5}$ $4.6\times 10^{-5}$ $\mathbf{12.4\times}$
9600 (late) $\mathbf{8.1\times 10^{-3}}$ $1.8\times 10^{-3}$ $7.0\times 10^{-4}$ $4.8\times$
Table 7: Rayleigh quotients and anisotropy (seed 271).
Step $q(\mathbf{v}_{\mathrm{b}})$ $q(\mathbf{v}_{\mathrm{sw}})$ $q(\mathbf{v}_{\mathrm{PC2}})$ Anisotropy $\alpha$
200 (init) $2.0\times 10^{-6}$ $1.1\times 10^{-6}$ $1.4\times 10^{-6}$ $1.6\times$
2200 (peak) $4\times 10^{-5} 1.1$$1\times 10^{-5}$ $2.5\times 10^{-5}$ $5.9\times$
4400 (transition) $9.5\times 10^{-5}$ $5.1\times 10^{-5}$ $5.2\times 10^{-5}$ $\mathbf{40.8\times}$
9200 (late) $1.0\times 10^{-4}$ $4.2\times 10^{-5}$ $7.6\times 10^{-5}$ $8.3\times$
Three observations emerge from Tables˜6 and 7:
  1. 1.
    The backbone stiffens progressively. $q(\mathbf{v}_{\mathrm{b}})$ increases by three orders of magnitude from initialization to late training (seed 42: $2.4\times 10^{-6}\to 8.1\times 10^{-3}$), outpacing all other directions.
  2. 2.
    Anisotropy spikes at the $\lambda$-transition. The moment when the probe loss weight doubles ($\lambda:2\to 4$ at step 4000) creates a sudden curvature increase along the backbone: $\alpha=12.4\times$ in seed 42, $40.8\times$ in seed 271. This is the sharpest curvature event in training.
  3. 3.
    The backbone is not the top Fisher eigenvector. Extracting the leading Fisher eigenvector $\mathbf{u}_{1}$ via the $M\times M$ Gram matrix trick
    $(\mathbf{G}\mathbf{G}^{\!\top})\,\mathbf{a}\;=\;\sigma^{2}\,\mathbf{a}\;\;\Longrightarrow\;\;\mathbf{u}_{1}\;=\;\mathbf{G}^{\!\top}\mathbf{a}_{1}\;/\;\lVert\mathbf{G}^{\!\top}\mathbf{a}_{1}\rVert,$ (29)
    the overlap $\lvert\langle\mathbf{u}_{1},\,\mathbf{v}_{\mathrm{b}}\rangle\rvert\approx 0.001$—essentially zero. The backbone captures cumulative drift, not the direction of steepest instantaneous curvature. It is a slow, persistent mode of the optimizer dynamics rather than a mode of the loss landscape Hessian.
@@ 4.5 Fisher Curvature: The Backbone Stiffens
Figure 3: Fisher curvature along the backbone stiffens over training (seed 42). Left: Rayleigh quotients $q(\mathbf{v})$ for the backbone (red), switch direction (blue), and random orthogonal directions (grey). The backbone curvature increases by three orders of magnitude. Right: Anisotropy ratio $\alpha=q(\mathbf{v}_{\mathrm{b}})/\operatorname{\mathbb{E}}[q(\mathbf{w}_{\perp})]$ spikes at the $\lambda$-transition (step 4000) before partially relaxing.
@@ 5.1 SGD-Family Control Experiment
We trained the same model under four optimizer configurations, using identical data, initialization (seed 42), schedule, and checkpointing protocol (Table˜8). The SGD variants differ from AdamW only in the optimizer; in particular, they lack per-parameter adaptive scaling.
@@
SGD without momentum failed to train (val loss remained >8>8 and probe accuracy stayed at chance). Both momentum-SGD variants (C and C) trained slowly but reached val $\approx$ 5.1, with weak probe OOD signal ($p_{\mathrm{ood}}\leq 0.015$) and no pronounced oscillatory switching. Nesterov momentum and decoupled weight decay produced negligible improvements over standard momentum-SGD at matched step: at step 2000, val loss differed by < 0.03{<}\,0.03 and probe OOD accuracy by ${\sim}0.001$ (Table˜9).
@@
These results indicate that momentum alone does not produce the multi-dimensional slow–fast structure observed under AdamW, and that the key ingredient is AdamW’s adaptive per-parameter scaling (Section˜4.2).
@@ 5.2 Matched-Loss Geometry
Even when AdamW passes through the same loss range early in training, its drift is already non-colinear: PC1 explains only 69–82% of row-normalized displacement energy with $k_{95}=2$33 over windows spanning this regime. In contrast, momentum-SGD remains nearly colinear at matched loss ($\rho_{1}\approx 0.98$$1.00$, $k_{95}=1$) and exhibits extremely small drift ($\lVert\Delta\bm{\theta}\rVert<1$) over the corresponding plateau windows. Decoupled weight decay (SGDW) and Nesterov momentum produce negligible changes relative to standard momentum-SGD (Table˜9).
@@ 6.2 Results
At the optimal learning rate ($6\times 10^{-4}$), the model reaches $p_{\mathrm{ood}}=0.782$—exceeding the training-time peak of 0.777—within 1000 reheating steps (Table˜11). The probe basin remains geometrically present in the late-training landscape. Reheating reveals that it is not erased but becomes dynamically suppressed by backbone stiffening.
@@ 6.3 Connection to Backbone Stiffening
The transient nature of reheating is consistent with the backbone stiffening documented in Section˜4.5. During training, the backbone direction accumulates Fisher curvature. By step 10,000, $q(\mathbf{v}_{\mathrm{b}})$ has increased by three orders of magnitude relative to initialization.
Reheating exposes the competition between transverse probe drive and backbone curvature. The backbone coordinate behaves as a stiffened slow variable, while transverse modes respond rapidly to increased probe weighting. Let the component of the probe gradient along the backbone be $g_{\parallel}(t)$ and the curvature along the backbone be $\kappa(t)\approx q(\mathbf{v}_{\mathrm{b}})$. The backbone coordinate evolves approximately as
$\frac{da}{dt}\;\approx\;-\eta_{t}\,g_{\parallel}(t)\;-\;\eta_{t}\,\kappa(t)\,a(t),$ (30)
where the first term is the probe-driven drift and the second is the curvature-mediated restoring force. When $\kappa$ is large (stiffened backbone), only a large $\eta_{t}$ can overcome it; as the cosine schedule reduces $\eta_{t}$, the restoring force dominates and the model returns to the LM basin.
The sign flip in update–backbone alignment (Section˜4.3) provides a complementary perspective. In early training, the optimizer drifts in the $\mathbf{v}_{\mathrm{b}}$ direction (toward the probe basin). After the sign flip ($t^{*}\approx 5000$), the drift reverses—the optimizer moves along $\mathbf{v}_{\mathrm{b}}$, away from the probe basin. Reheating temporarily reverses this via the strong probe gradient, but once $\eta_{t}$ decays, the backbone’s natural $\mathbf{v}_{\mathrm{b}}$ drift reasserts itself.
@@ 8 Discussion
8.1 Slow–Fast Decomposition of Optimizer Dynamics
The backbone–transverse decomposition reveals a structural separation in transformer training under AdamW. Long-horizon parameter evolution concentrates in a low-dimensional drift direction, while oscillatory objective competition unfolds in a high-dimensional transverse subspace.
This separation is not imposed by architecture or by explicit constraints in the loss. It emerges from optimizer dynamics. Momentum integrates weak but temporally coherent gradient bias, and adaptive normalization suppresses incoherent variance. The result is an effective timescale separation: a slow cumulative drift manifold and faster transverse fluctuations.
Formally, the dynamics resemble a singularly perturbed system,
$\frac{d\bm{\theta}}{dt}\;=\;f_{\mathrm{slow}}(\bm{\theta})\;+\;\epsilon^{-1}\,f_{\mathrm{fast}}(\bm{\theta}),$ (34)
where the slow component corresponds to backbone drift and the fast component to transverse switching. The scale separation parameter reflects the ratio between temporal gradient coherence and transverse fluctuation variance. This framing shifts the focus of optimization analysis from static curvature to trajectory-level structure.
8.2 Instantaneous Freedom vs. Cumulative Coherence
The backbone does not imply that gradients are low-dimensional. At each step, gradients remain high-dimensional and largely isotropic relative to the backbone. The distinction lies between instantaneous variability and cumulative displacement.
In high dimensions, large transverse components can dominate the norm of individual updates yet cancel over time if their directions fluctuate. A much smaller but temporally coherent component can therefore dominate long-horizon displacement. The backbone reflects this integrated coherence rather than a restriction of gradient expressivity.
This distinction clarifies why the backbone is nearly orthogonal to top Fisher eigenvectors. It is not the steepest local direction of the loss landscape, but the direction in which optimizer integration accumulates persistent bias.
8.3 Optimizer-Induced Geometry
The SGD-family controls demonstrate that this structure is not a generic property of the objective. Under SGD with or without momentum, trajectories remain nearly colinear and fail to develop multi-dimensional drift structure, even at matched validation loss.
Adaptive per-parameter scaling appears to be the key ingredient. By normalizing coordinates according to gradient variance, AdamW selectively amplifies temporally coherent components and suppresses incoherent ones. This produces qualitatively richer cumulative geometry than momentum alone.
These results suggest that optimizer choice influences not only convergence speed or stability, but the geometric structure of the training trajectory itself.
8.4 Dynamical Consequences: Reheating and Basin Accessibility
Reheating experiments provide a dynamical probe of this structure. Transverse probe dynamics can be transiently re-excited from late-training checkpoints, but accumulated backbone drift remains largely intact. As curvature along the backbone increases, larger learning rates are required to overcome the associated restoring force. When the learning rate decays, the system relaxes back toward the backbone-aligned regime.
Thus reheating does not contradict backbone dominance; it reveals that switching dynamics correspond to transverse excursions around a persistent slow manifold shaped by optimizer integration.
8.5 Limitations and Scope

### v2

@@ 4.9 Backbone Stiffening Under Rotation
Thus backbone rotation and stiffening are coupled: they are two aspects of the same dynamical reorganization triggered by the objective reweighting.
@@
Figure 3: Fisher curvature along key directions (seed 42). Top: Rayleigh quotient $q(\mathbf{v})$ for the backbone (blue), switch direction (red), second PC (green), and mean over random orthogonal directions (grey). The backbone curvature increases by three orders of magnitude. Bottom: Anisotropy ratio $\alpha=q(\mathbf{v}_{\mathrm{b}})/\operatorname{\mathbb{E}}[q(\mathbf{w}_{\perp})]$ spikes at the $\lambda$-transition (step 4000) before partially relaxing.
4.10 Geometric Interpretation
The cumulative trajectory admits the decomposition
$\theta(t)=\theta(0)+a(t)\,v_{b}(t)+r(t),\qquad r(t)\perp v_{b}(t),$
where $a(t)$ tracks slow drift along the manifold and $r(t)$ captures transverse dynamics.
Because adjacent rolling tangents maintain high alignment ($\rho>0.7$), the manifold has bounded local curvature. But because phase-level backbones satisfy $|\langle v_{E},v_{L}\rangle|\approx 0.32$, the manifold is globally curved by $\approx 71^{\circ}$.
The $\lambda$-transition concentrates curvature in the 4k–6k region, producing a transient geometric dead zone (both $A_{E}$ and $A_{L}$ small) through which the slow direction rotates. The post-switch phase is more rank-1 concentrated (PC1 $\approx 81\%$) but evolves in a nearly orthogonal direction.
4.11 Summary of Mechanism
The backbone is:
  • not aligned with instantaneous gradients,
  • not the top Fisher eigenvector,
  • not a static direction fixed by the loss landscape.
It is instead:
  • a locally dominant tangent to a curved slow manifold,
  • induced by optimizer integration of temporally coherent gradient bias,
  • stabilized by second-moment normalization ($\beta_{2}$ ablation: PC1 drops 68% $\to$ 52% at $\beta_{2}=0$),
  • reoriented by objective reweighting ($|\langle v_{E},v_{L}\rangle|\approx 0.32$ at the $\lambda$-switch),
  • and correlated with task performance: $\mathrm{corr}(p_{\mathrm{ood}},\|r\|)\approx+0.8$ early, $\approx-0.8$ late.
The optimizer does not merely change convergence speed — it shapes the geometry of the cumulative trajectory, creating a smooth slow manifold whose curvature is controlled by the training objective.
@@ 5.1 SGD-Family Control Experiment
We trained the same model under four optimizer configurations, using identical data, initialization (seed 42), schedule, and checkpointing protocol (Table˜3). The SGD variants differ from AdamW only in the optimizer; in particular, they lack per-parameter adaptive scaling.
@@
SGD without momentum failed to train (val loss remained >8>8 and probe accuracy stayed at chance). Both momentum-SGD variants (C and C) trained slowly but reached val $\approx$ 5.1, with weak probe OOD signal ($p_{\mathrm{ood}}\leq 0.015$) and no pronounced oscillatory switching. Nesterov momentum and decoupled weight decay produced negligible improvements over standard momentum-SGD at matched step: at step 2000, val loss differed by < 0.03{<}\,0.03 and probe OOD accuracy by ${\sim}0.001$ (Table˜4).
@@
These results indicate that momentum alone does not produce the multi-dimensional slow–fast structure observed under AdamW, and that the key ingredient is AdamW’s adaptive per-parameter scaling (Section˜4.7).
@@ 5.2 Matched-Loss Geometry
Even when AdamW passes through the same loss range early in training, its drift is already non-colinear: PC1 explains only 69–82% of row-normalized displacement energy with $k_{95}=2$33 over windows spanning this regime. In contrast, momentum-SGD remains nearly colinear at matched loss ($\rho_{1}\approx 0.98$$1.00$, $k_{95}=1$) and exhibits extremely small drift ($\lVert\Delta\bm{\theta}\rVert<1$) over the corresponding plateau windows. Decoupled weight decay (SGDW) and Nesterov momentum produce negligible changes relative to standard momentum-SGD (Table˜4).
@@
6 Effect of $\beta_{2}$ on Cumulative Drift Structure
To test whether backbone dominance depends on second-moment normalization strength, we vary $\beta_{2}$ in AdamW while keeping architecture, data, learning-rate schedule, and objective weighting fixed. All runs are trained for 4,000 steps with $\lambda=2.0$.
6.1 Backbone dominance degrades smoothly with lower $\beta_{2}$
For each $\beta_{2}\in\{0.99,0.95,0.90,0.80\}$, we compute the uncentered trajectory PCA over the trunk parameters (D=25,165,824D=25{,}165{,}824). The fraction of cumulative displacement explained by the first principal component decreases monotonically as $\beta_{2}$ decreases:
$\beta_{2}$ PC1% (row-norm) PC1% (raw) Drift magnitude Best $p_{\mathrm{ood}}$
0.99 68.8 83.1 120 0.951
0.95 68.3 82.5 123 0.939
0.90 66.7 80.8 128 0.814
0.80 63.8 79.0 141 0.682
0.0 52.5 81.3 $1.6\times 10^{5}$ 0.005
At $\beta_{2}=0.0$ (no second-moment normalization), drift magnitude explodes by ${\sim}1{,}300\times$ (from 120 to $1.6\times 10^{5}$), PC1 concentration drops from 69% to 53%, and probe accuracy collapses entirely. Lower $\beta_{2}$ therefore weakens low-dimensional dominance without altering model architecture or task.
6.2 Power-law regimes persist across $\beta_{2}$
Despite quantitative changes in backbone dominance, the qualitative two-phase structure remains consistent across $\beta_{2}$. We fit $|a(t)|=C_{a}\,t^{\gamma_{a}}$ and $\|r(t)\|=C_{r}\,t^{\gamma_{r}}$ in two windows:
$\gamma_{a}$ (backbone) $\gamma_{r}$ (residual)
$\beta_{2}$ 0–2k 2k–4k 0–4k 0–2k 2k–4k 0–4k
0.99 +1.97+1.97 0.01-0.01 +1.32+1.32 +0.78+0.78 +0.36+0.36 +0.55+0.55
0.95 +1.94+1.94 0.03-0.03 +1.31+1.31 +0.78+0.78 +0.35+0.35 +0.57+0.57
0.90 +1.93+1.93 0.03-0.03 +1.31+1.31 +0.84+0.84 +0.32+0.32 +0.60+0.60
0.80 +1.98+1.98 0.05-0.05 +1.35+1.35 +0.91+0.91 +0.21+0.21 +0.62+0.62
0.0 +2.45+2.45 0.36-0.36 +1.84+1.84 +1.08+1.08 0.44-0.44 +0.56+0.56
All early-phase fits have $R^{2}>0.98$ for $\gamma_{a}$ and >0.88>0.88 for $\gamma_{r}$, confirming superlinear backbone growth ($\gamma_{a}\approx 1.9$$2.0$) independent of $\beta_{2}$. In the later phase, the backbone saturates ($\gamma_{a}\approx 0$) for $\beta_{2}\geq 0.80$, while at $\beta_{2}=0.0$ the backbone actively retreats ($\gamma_{a}=-0.36$) and the residual contracts ($\gamma_{r}=-0.44$).
Thus $\beta_{2}$ modulates the magnitude of late-phase dynamics rather than changing the qualitative trajectory regime.
6.3 Update alignment and objective robustness
We next measure the signed alignment between accumulated 200-step updates and the rolling backbone direction:
$\beta_{2}$ Mean $|\cos(u,v_{b})|$ Early (<<2k) $|\cos|$ $\mathrm{corr}(p_{\mathrm{ood}},\|r\|)$ early
0.99 0.243 0.368 +0.86+0.86
0.95 0.244 0.363 +0.85+0.85
0.90 0.234 0.343 +0.40+0.40
0.80 0.227 0.328 +0.39+0.39
0.0 0.238 0.269 0.57-0.57
Peak early alignment decreases gradually with lower $\beta_{2}$ ($0.37\to 0.27$), indicating reduced temporal coherence of cumulative updates. Probe robustness follows the same monotonic trend: the best out-of-distribution probe accuracy decreases smoothly as $\beta_{2}$ decreases.
The early-phase correlation $\mathrm{corr}(p_{\mathrm{ood}},\|r\|)$ reveals a qualitative shift: for $\beta_{2}\geq 0.95$, residual growth accompanies improving probe accuracy ($r\approx+0.85$); for $\beta_{2}=0.0$, the correlation is negative (0.57-0.57), indicating that without second-moment normalization, transverse dynamics are destructive rather than productive.
Together, these observations indicate that second-moment normalization strength controls the coherence of accumulated drift rather than the instantaneous gradient geometry. Larger $\beta_{2}$ produces stronger directional persistence and greater low-dimensional dominance; smaller $\beta_{2}$ increases dispersion across directions and weakens probe stability.
6.4 Summary
Varying $\beta_{2}$ provides a graded control on backbone dominance. With all other factors held fixed, reducing second-moment normalization strength:
  • weakens low-dimensional trajectory concentration (PC1: $68.8\%\to 52.5\%$),
  • decreases early update–backbone alignment ($0.37\to 0.27$),
  • increases drift magnitude ($120\to 1.6\times 10^{5}$),
  • reduces probe robustness ($p_{\mathrm{ood}}:0.95\to 0.005$),
  • and reverses the sign of early residual–performance correlation ($0.86\to-0.57$).
These results support the interpretation that backbone structure arises from optimizer-integrated update dynamics rather than architectural constraints alone.
@@ 7.2 Results
At the optimal learning rate ($6\times 10^{-4}$), the model reaches $p_{\mathrm{ood}}=0.782$—exceeding the training-time peak of 0.777—within 1000 reheating steps (Table˜6). The probe basin remains geometrically present in the late-training landscape. Reheating reveals that it is not erased but becomes dynamically suppressed by backbone stiffening.
@@ 7.3 Connection to Backbone Stiffening
The transient nature of reheating correlates with increased Fisher curvature along the backbone direction, documented in Section˜4.9. During training, $q(\mathbf{v}_{\mathrm{b}})$ increases by three orders of magnitude from initialization to step 10,000.
When curvature along the backbone is large, larger learning rates are required to perturb accumulated drift; as the learning rate decays, the system returns toward the LM-dominant regime. This is consistent with a picture in which transverse modes respond rapidly to increased probe weighting, while the backbone coordinate resists perturbation proportionally to accumulated curvature.
The sign flip in update–backbone alignment (Section˜4.8) provides a complementary observation. In early training, the optimizer drifts in the $\mathbf{v}_{\mathrm{b}}$ direction (toward the probe basin). After the sign flip ($t^{*}\approx 5000$), the drift reverses—the optimizer moves along $\mathbf{v}_{\mathrm{b}}$, away from the probe basin. Reheating temporarily reverses this via the strong probe gradient, but once $\eta_{t}$ decays, the accumulated drift direction reasserts itself.
@@ 9 Discussion
9.1 Empirical Decomposition of Training Dynamics
We presented an empirical decomposition of transformer training under AdamW into two geometric components:
  1. 1.
    A dominant cumulative drift direction (the rolling backbone), defined operationally via uncentered trajectory PCA in rolling windows.
  2. 2.
    Transverse residual dynamics associated with probe switching.
This decomposition is descriptive rather than assumed. It follows directly from measured drift geometry. Across two seeds and 10,000 training steps, the backbone is:
  • Low-dimensional: PC1 captures 71–72% of row-normalized variance over the full trajectory, and 60–81% in phase-specific windows (§4.1, §4.3).
  • Locally smooth: Adjacent-window cosine $\rho(t)>0.7$ everywhere, with mean $\approx 0.80$ (§4.2).
  • Globally reoriented: Early–late phase cosine $|\langle v_{E},v_{L}\rangle|=0.32$ ($\approx 71^{\circ}$ rotation), with curvature concentrated in the 4k–6k transition region (§4.3–4.4).
The quantitative consistency across seeds (all reported values agree to within 1–2% between seed 42 and seed 271) indicates that these geometric features are robust properties of the training dynamics rather than seed-specific artifacts.
9.2 Instantaneous Gradients vs Accumulated Updates
[…]
We do not claim a formal dynamical-systems separation. We only observe that accumulated update geometry differs qualitatively from per-step gradient geometry, and that this difference is controlled by optimizer configuration (§4.7).
9.3 Optimizer Dependence
Optimizer ablations provide causal support for the claim that backbone geometry is optimizer-induced:
[…]
  • SGD control: SGD-family variants fail to reproduce the same learning outcome. AdamW with the same architecture achieves PC1 $\approx 81\%$ on the same trajectory interval.
  • Monotonic degradation: Across $\beta_{2}\in\{0.99,0.95,0.90,0.80,0.0\}$, backbone concentration, update alignment, and best probe accuracy all degrade monotonically.
These effects occur under matched architecture, data, and schedules, isolating optimizer configuration as the primary variable. We therefore conclude that backbone geometry is optimizer-induced rather than a generic property of the loss landscape.
9.4 Phase Reorientation Under Objective Reweighting
The $\lambda$-switch at step 4000 ($\lambda:2\to 4$) coincides with a measurable dynamical phase transition:
  • Backbone rotation: The slow direction reorients by $\approx 71^{\circ}$, passing through a geometric dead zone ($A_{E}\approx 0.02$, $A_{L}\approx 0.16$) near step 3500 (§4.4).
  • Power-law regime change: Backbone growth exponent drops from $\gamma_{a}=+1.74$ (pre-switch) to 0.08-0.08 (post-switch); residual exponent drops from $\gamma_{r}=+0.84$ to 0.31-0.31 (§4.5).
  • Sign reversal: Update–backbone alignment flips from $\cos\approx-0.32$ to +0.11+0.11 near step 5100 (§4.8).
[…]
  • Correlation flip: $\mathrm{corr}(p_{\mathrm{ood}},\|r\|)$ reverses from +0.85+0.85 (0–4k) to 0.76-0.76 (6k–10k) (§4.6).
We do not claim that curvature causes rotation. Rather, we document that objective reweighting produces correlated, measurable changes across five independent geometric diagnostics. The rolling tangent remains locally smooth throughout ($\rho>0.7$); the phase transition reflects gradual global reorientation concentrated in the 4k–6k region.
9.5 Reheating as a Geometric Probe
Reheating experiments reveal that:
  • Probe performance can be transiently restored from late checkpoints.
  • Re-entry magnitude depends on learning rate and $\beta_{2}$.
  • Accumulated backbone drift remains largely unchanged during reheating.
These findings are consistent with the backbone–residual decomposition: reheating perturbs transverse components while leaving dominant cumulative drift largely intact. The correlation flip documented in §4.6 provides additional context: after step 6000, residual re-growth is associated with declining probe accuracy ($r\approx-0.76$0.79-0.79), suggesting that late residual dynamics drive the model away from the probe solution rather than toward it.
9.6 The Three Dynamical Phases
The power-law analysis (§4.5) reveals that training decomposes into three phases with distinct geometric signatures:
  1. 1.
    Acceleration (0–2k): Both backbone and residual grow rapidly ($a\sim t^{2.1}$, $r\sim t^{1.2}$). The trajectory explores broadly, with productive residual energy ($\mathrm{corr}(p_{\mathrm{ood}},\|r\|)\approx+0.8$).
  2. 2.
    Consolidation (2k–6k): Backbone growth slows ($\gamma_{a}:2.1\to 0.67\to 0.15$) while the residual contracts ($\gamma_{r}\approx-0.4$ to 1.4-1.4). The $\lambda$-switch at step 4000 sharpens the contraction. The model consolidates along the slow direction.
  3. 3.
    Plateau (6k–10k): The backbone retreats ($\gamma_{a}\approx-0.19$), the residual re-grows ($\gamma_{r}\approx+0.34$), and probe accuracy declines. The drift direction becomes orthogonal to both phase backbones ($A_{E}\approx A_{L}\approx 0.02$$0.18$). The late-phase rank-1 concentration (PC1 $\approx 81\%$) indicates tight low-rank drift, but in a direction no longer aligned with either learning phase.
These three phases are not assumed a priori; they emerge from piecewise power-law fits with high $R^{2}$ values (>0.85>0.85, most >0.97>0.97) and are consistent across both seeds.
9.7 Scope and Limitations